OpenStudy (xcecex):

SOMEONE PLEASE HALP MEEEEEEEEE!!! ill medal and fan!!! Meg hired Rich to help her at a car wash. Meg earned $5.40 per car they washed, plus $5 in tips. Rich earned $3.20 per car they washed, plus $8 in tips. Let c represent the number of cars Meg and Rich washed. Which simplified expression represents the total amount of money they earned? A. 21.6c B. 13.4c + 8.2 C. 8.6c + 13 D. 10.4c + 11.2c

OpenStudy (xcecex):

This is probably baby math to you but i just need help on how to do it/ break it down so its simpler :l

Directrix (directrix):

No problem is baby math to me. A problem is hard if I cannot do it. A problem is easy if I can do it.

OpenStudy (xcecex):

Oh, well then >_>

Directrix (directrix):

Assumption: Meg and Ritch worked together on the same cars. Meg: 5.40 per car --> income is 5.40 * c + 5 Ritch: 3.20 per car --> income is 3.20 * c + 8

OpenStudy (xcecex):

Right, so then now that i know their incomes what do i do next?

Directrix (directrix):

Which simplified expression represents the total amount of money they earned? M + R = 5.40 * c + 5 + 3.20 * c + 8 = ? @XceceX

Directrix (directrix):

5.40 * c + 3.20 * c + 8 + 5 =

Directrix (directrix):

Post what you get, okay?

OpenStudy (xcecex):

Okay :)

OpenStudy (xcecex):

21.6c? @Directrix

Directrix (directrix):

No.

Directrix (directrix):

5.40 * c + 3.20 * c = 8.60 c 8 + 5 = 13 8.60 c + 13 = which option. --> look above 8.60c and 13 are not like terms.

Directrix (directrix):

@XceceX Which option?

OpenStudy (xcecex):

A srry i went to go eat @Directrix